# Direct Preference Knowledge Distillation for Large Language Models

### v1

@@ 4.4 Variants of Preference Objective
In the subject of fine-tuning language models with human feedback, there are many optimization methods based on preference. These different optimization methods are all based on the original work DPO, and modify the preference form, normalization method, etc. on this basis. In the field of knowledge distillation of large models, the function form derived in this paper is a relatively basic form. So similarly, we also conducted preliminary experiments on different forms of preference calculation methods, and hope to bring inspiration to subsequent research. The experiments conducted in this chapter did not perform special parameter fine-tuning.
@@
Table 2: Variants of preference objective. The definition of each preference is as follows: (1) SimPO = $-\log\sigma\left(\frac{\beta}{\left|y_{t}\right|}\log q_{\theta}\left(y_{t}|x\right)-\frac{\beta}{\left|y_{s}\right|}\log q_{\theta}\left(y_{s}|x\right)-\gamma\right)$;
(2) CPO = $-\log\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}|x\right)}{q_{\theta}\left(y_{t}|x\right)}\right)-\log q_{\theta}\left(y_{t}|x\right)$;
(3) IPO = $\left(\log\frac{q_{\theta}\left(y_{t}|x\right)}{p\left(y_{t}|x\right)}-\log\frac{q_{\theta}\left(y_{s}|x\right)}{p\left(y_{s}|x\right)}-\frac{1}{2\tau}\right)^{2}$.
We experimented with four forms of preference: IPO, CPO, SimPO. Their definitions and experimental results are shown in Figure 2.
From the experimental results, we could see that although other forms of preference do not exceed our method, they also show certain effectiveness. Pal et al. (2024) showed that in the context of learning from preference data, the form of the preference also affects how the relative scores of $y_{t}$ and $y_{s}$ grow during the training process. Our results show that these problems and solutions are also applicable to the field of knowledge distillation, which also provides inspiration for future work.
@@ 4.5 Ablation Studies
We conduct ablation experiments on the experimental setting and measure the scores on the GPT-2 Base model and DollyEval. We evaluate the results of removing length normalization and language modeling loss. In Table 3, KLD and rKLD represent the minimum values of KL divergence and reverse KL divergence from the teacher model during training, respectively. To ensure consistent distribution conditions, the distribution difference is calculated based on the first token output. The reward is the maximum value during training and is calculated in the same way as above. From the results, we can see that the components added in the experiment have a positive effect on the results, especially the improvement of implicit reward. The complete method performs well with the teacher model under both reverse KL and forward KL metrics, and has the highest implicit reward and score.
@@ 4.6 Results of Different Generation Lengths
We experiment on subsets of the test set with different response lengths to test the effectiveness of our approach for scenarios with different output length requirements. We divide the DollyEval test set into three subsets according to the length of the golden response: $\left(0,30\right)$, $\left(30,70\right)$ and $\left(70,\infty\right)$, where 30 and 70 are the median and mean of the response length, respectively. We then evaluate the response results of different baselines and DPKD methods on these subsets.
The results of RougeL and exact match are shown in Figure 4 and Table 4. Our DPKD method performs well on all subsets of response length splits. In particular, DPKD exceeds the baseline by the most in the results with response lengths in the middle. In the exact match results, the scores of many baseline methods drop to 0 when the golden response length is longer, while DPKD still performs well and exceeds the baseline.
@@ 5 Related Work
is a widely used technique in model training and fine-tuning, and was first proposed in the work Bucila et al. (2006), Hinton et al. (2015) further enriched and conducted studies on KD. Researchers have explored the application of KD in text generation tasks Song et al. (2020); Zhang et al. (2023); Jiao et al. (2020); Sun et al. (2019). The standard form of KD in NLP is to minimize the KLD between student model and teacher, including train on teacher-generated text Kim and Rush (2016); Taori et al. (2023) and teacher output at each step Sanh et al. (2019). There are also many works Agarwal et al. (2023); Passalis and Tefas (2018); Wen et al. (2023) exploring the metric of distribution distance in KD. Work Huang et al. (2022) points out the inadequacy of KL distance when the difference between teacher and student distribution is large, and proposed a relaxed matching term as a supplement to KL divergence. Concurrent works Cho and Hariharan (2019); Mirzadeh et al. (2020) also proposed a similar phenomenon when the student and teacher model sizes significantly differ.
@@
has received attention in recent years due to the booming open source LLMs community Zhang et al. (2022a); Touvron et al. (2023). KD in large models is usually divided into black-box distillation Taori et al. (2023); Chiang et al. (2023) and white-box distillation Kim et al. (2024). The black-box KD takes the text generated by the teacher as knowledge and performs distillation, while the wite-box KD use the model intermediate layers or output logits distribution Jiao et al. (2020); Wang et al. (2020). Recent works have recognized the inadequacy of KL divergence in knowledge distillation of LLMs Gu et al. (2023) and applied reverse KL divergence to KD of LLMs. Concurrent work Wu et al. (2024) illustrates the shortcomings of reverse KLD and KLD and designs a KL divergence that adaptively allocates weights.
@@ Appendix A Complete Derivation of DPKD Objective
where $q^{*}(y|x)\mathrel{\coloneqq}\frac{1}{Z(x)}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$, and $Z\left(x\right)$ is the normalized function of the distribution, which is required to be independent of $\theta$ and $y$. It is defined as $Z(x)\mathrel{\coloneqq}\sum_{y}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$.
@@
Since $Z\left(x\right)$ is a normalized function independent of $\theta$, the minimum value from the optimization objective  19 is achieved by minimizing the first KL divergence term. We can derive the optimal solution of Equation 19:

### v2

@@ 4.4 Variants of Preference Objective
In the subject of fine-tuning language models with human feedback, there are many optimization methods based on preference. These various optimization techniques are all based on the original work DPO, and modify aspects such as the preference format and normalization methods. In the context of knowledge distillation for large models, the function form derived in this paper represents a relatively basic form. So similarly, we also conducted preliminary experiments on different methods for calculating preferences, with the aim of inspiring future research. It is important to note that the experiments conducted in this chapter did not perform special parameter fine-tuning.
@@
Table 2: Variants of preference objective. The definitions of each preference are as follows: (1) SimPO = $-\mathbb{E}\log\sigma\left(\frac{\beta}{\left|y_{t}\right|}\log q_{\theta}\left(y_{t}|x\right)-\frac{\beta}{\left|y_{s}\right|}\log q_{\theta}\left(y_{s}|x\right)-\gamma\right)$;
(2) CPO = $-\mathbb{E}\left[\log\sigma\left(\beta\log\frac{q_{\theta}\left(y_{t}|x\right)}{q_{\theta}\left(y_{t}|x\right)}\right)-\log q_{\theta}\left(y_{t}|x\right)\right]$;
(3) IPO = $\mathbb{E}\left(\log\frac{q_{\theta}\left(y_{t}|x\right)}{p\left(y_{t}|x\right)}-\log\frac{q_{\theta}\left(y_{s}|x\right)}{p\left(y_{s}|x\right)}-\frac{1}{2\tau}\right)^{2}$.
We experimented with three other forms of preference: IPO, CPO, SimPO. Their experiment results and definitions are shown in Table 2. From the experimental results, we could see that although other forms of preference do not surpass our method, they demonstrate a certain level of effectiveness. Pal et al. (2024) illustrated that in the context of learning from preference data, the form of the preference also affects how the relative scores of $y_{t}$ and $y_{s}$ grow during the training process. Our findings suggest that these problems and solutions are also applicable to the field of knowledge distillation.
@@ 4.5 Ablation Studies
We conduct ablation experiments on the experimental settings and measure the scores on the GPT-2 Base model and DollyEval. We evaluate the results of omitting length normalization and language modeling loss. In Table 3, KLD and rKLD represent the minimum values of KL divergence and reverse KL divergence from the teacher model during training, respectively. To ensure consistent distribution conditions, the distribution difference is calculated based on the output of the first token. The reward is the maximum value during training and is calculated using the same method described above. From the results, we can see that the components introduced in the experiment have a positive effect on the results, especially the improvement of implicit reward. The comprehensive method performs well with the teacher model under both reverse KL and forward KL metrics, achieving the highest implicit reward and score.
@@ 4.6 Results of Different Generation Lengths
In this section, we conduct experiments on subsets of the test set with varying response lengths to assess the effectiveness of our approach in scenarios with different output length requirements. We divide the DollyEval test set into three subsets according to the length of the golden response: $\left(0,30\right)$, $\left(30,70\right)$ and $\left(70,\infty\right)$, where 30 and 70 are the median and mean of the response lengths, respectively. We then evaluate the response results of various baseline methods and DPKD on these subsets.
The results of the RougeL and exact match are shown in Figure 5 and Table 4. Our DPKD method performs well on all subsets of response length splits. Notably, DPKD exceeds the baseline by the most in the results with response lengths in the middle. In the exact match results, the scores of many baseline methods drop to 0 when the golden response length is longer, while DPKD still performs well and exceeds the baseline.
@@ 5 Related Work
is a widely used technique in model training and fine-tuning, and was first proposed in the work Bucila et al. (2006), Hinton et al. (2015) further enriched and conducted studies on KD. Researchers have explored the application of KD in text generation tasks Song et al. (2020); Zhang et al. (2023); Jiao et al. (2020); Sun et al. (2019). The standard form of KD in NLP is to minimize the KLD between student model and teacher, including train on teacher-generated text Kim and Rush (2016); Taori et al. (2023) and teacher output at each step Sanh et al. (2019). There are also many works Agarwal et al. (2023); Passalis and Tefas (2018); Wen et al. (2023) exploring the metric of distribution distance in KD. Works Huang et al. (2022); Cho and Hariharan (2019); Mirzadeh et al. (2020) proposed inadequacy of KLD when the student and teacher model sizes significantly differ.
@@
has received attention in recent years Zhang et al. (2022a); Touvron et al. (2023). KD in large models is usually divided into black-box distillation Taori et al. (2023); Chiang et al. (2023) and white-box distillation Kim et al. (2024). The black-box KD takes the text generated by the teacher as knowledge and performs distillation, while the wite-box KD use the model intermediate layers or output logits distribution Jiao et al. (2020); Wang et al. (2020). Recent works have recognized the inadequacy of KL divergence in knowledge distillation of LLMs Gu et al. (2023) and applied reverse KL divergence to KD of LLMs. Concurrent work Wu et al. (2024) illustrates the shortcomings of reverse KLD and KLD and designs a KL divergence that adaptively allocates weights.
@@ Appendix A Complete Derivation of DPKD Objective
where $q^{*}(y|x)\mathrel{\coloneqq}\frac{1}{Z(x)}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$, and $Z\left(x\right)$ is the scaling function of the distribution, which is required to be independent of $\theta$ and $y$. It is defined as $Z(x)\mathrel{\coloneqq}\sum_{y}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$.
@@
Since $Z\left(x\right)$ is a scaling function independent of $\theta$, the minimum value from the optimization objective  19 is achieved by minimizing the first KL divergence term. We can derive the optimal solution of Equation 19: